# Collaborative and parametric insurance on the Ethereum blockchain

### v2

@@ 1 Introduction
A blockchain is a decentralized ledger maintained by consensus within a peer-to-peer network. The nodes in this network agree on a shared history of data by executing an algorithm known as a consensus protocol. The Bitcoin blockchain, as described by Nakamoto [24], simply record transactions between bitcoin users. In contrast, the Ethereum blockchain, introduced by Buterin [6], is also equipped with a compiler to execute pieces of code called smart contract. Ethereum is designed to operate as a "world computer," enabling developers to deploy decentralized applications (dApps) that users can interact with directly.
Through these decentralized applications, Ethereum extends the capabilities of blockchain beyond simple fund transfers, enabling more complex financial operations and fostering the emergence of a new financial ecosystem known as decentralized finance (DeFi), as described by Schär [27]. Traditional finance relies on trusted third parties and centralized databases, which involve intermediary costs, lack transparency, operate only during limited hours, and are susceptible to cyberattacks. In contrast, distributed ledgers are publicly accessible, replicated in multiple locations, and function continuously, 24/7. Furthermore, the code that governs decentralized applications is auditable, improving transparency. For further discussion, we refer the reader to Lipton and Treccani [20, Chapter 2].
@@
In recent years, parametric insurance schemes have emerged in mass-market products, such as flight delay insurance offered by China Southern Airlines, which provides payouts of up to 300 yuan (approximately US\$40) for delays of three hours or more.
The main downside of parametric insurance is basis risk, where the insured’s losses may not match the coverage amount, or the parameter may not trigger. Effective structuring and pricing require understanding the policyholder’s exact risks and choosing the right parameter. The availability of finer granularity data and advances in predictive modeling can mitigate basis risk. Hybrid insurance solutions that combine parametric and standard indemnity mechanisms can also be effective, see for instance Lopez and Nkameni [21]. An initial payment following the parametric trigger can provide immediate relief, and this amount can later be deducted from the compensation resulting from the traditional claims assessment process. The policyholder has the option to either maintain parametric insurance or purchase an additional traditional insurance contract to cover basis risk. It is conceivable that some insurers might be willing to supplement parametric insurance in certain cases, as insurers and reinsurers often find themselves in this position. For instance, they might partially hedge their traditional insurance risk using a parametric Insurance-Linked Securities (ILS) solution, such as catastrophe bonds or mortality bonds. However, there has been a shift from parametric to indemnity triggers in the ILS sector, which means some challenges would need to be addressed. Note that death benefits in traditional insurance often consist of a fixed amount that varies according to limited number of parameters, such as the number of children or marital status. This forms a prominent example of a parametric insurance contract that is widely accepted and commercialized.
@@
We present a smart contract on the Ethereum blockchain to manage a portfolio of parametric insurance policies. We have developed an actuarial framework suited for weather-linked triggering events, specifically monitoring daily rainfall. This framework establishes ground rules for underwriting, rate-making, and managing risk through the determination of solvency capital. Our smart contract is designed to operate with minimal human interference, maximizing automation. Every interaction with the smart contract incurs a cost, as any transaction submitted to the blockchain requires a fee to be processed by the network of nodes. Therefore, optimizing the number of transactions is crucial for reducing operational costs. Actuarial methods for calculating risk capital are often too complex to be directly implemented in a high-level language like Solidity, which, for instance, only supports integer-based arithmetic. Additionally, deploying a smart contract on the blockchain entails costs that grow with its complexity, emphasizing the need for simplicity in design. One of our key contributions is the simplification of these computations through justified approximations. In particular, we propose a straightforward recursive formula to determine the solvency capital required to adequately cover the risks of a parametric insurance portfolio.
@@
The smart contract structure is inspired by the discussion in Cousaert et al. [10], from which we adopt the role distribution and the "tokenization" mechanism of risk. The smart contract owner writes the code and sets the initial values for parameters such as the premium loading and the risk level required for risk capital allocation. Investors participate by locking funds in the smart contract, in exchange for which they receive protocol tokens. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. The contribution of each investor is reflected in their token holdings. The value of the token, or equivalently its exchange rate against ETH, fluctuates over time depending on whether compensation are being paid to policyholders. Conversely, the token value increases if no compensation is paid at settlement. Token ownership enables governance of the insurance protocol. The contract owner may update the smart contract parameters based on recommendations from a board composed of token holders. This governance model transforms the smart contract into a Decentralized Autonomous Organization (DAO). The tokenization of risk on the blockchain can be viewed as a novel way to design insurance-linked securities, as outlined in Barrieu and Albertini [4]. By enabling policyholders to become investors, our smart contract creates a collaborative insurance scheme, akin to those described in Feng [12]. Investors, referred to as Surplus Providers (SPs), are analogous to liquidity providers in decentralized exchanges, which serve as blockchain-based trading venues for crypto assets; see, for instance, Mohan [23]. Pooling resources is not uncommon in the context of parametric insurance, as both the CCRIF and ARC are multi-country agreements to combat climate risk.
@@
This paper aims at showing the feasibility of deploying a beta version of such a smart contract, where certain complexities are intentionally simplified. We also identify the main limitations of this simplified framework and provide the research and actuarial communities with an open-source beta version of the smart contract. This prototype serves as a foundation for developing generalizations and improved versions that address these limitations.
The rest of the paper is organized as follows. Section 2 provides a brief overview on blockchain technology. Section 3 presents an actuarial framework for managing parametric insurance contract portfolios. Section 4 focuses on the design of the smart contract. We introduce the participants, including the contract owner, surplus providers, and policyholders, in Section 4.1. A smart contract is characterized by state variables, which we describe in Section 4.2. Examples of state variables include the smart contract’s balance and the tokens held by each participant. Users interact with the smart contract by invoking functions, and the possible actions and their impacts on the system’s state are detailed in Section 4.3. Section 5 demonstrates how the smart contract operates through an event-driven scenario, showcasing all its functions and culminating in the liquidation and reset of the smart contract. The main limitations of this beta version of the smart contract are discussed in Section 6, along with potential ideas for addressing them.
@@ 2 Blockchain technology and smart contracts
However, the Bitcoin blockchain was designed with a narrow scope: it enables peer-to-peer transfer of value but offers very limited programmability. As a result, more complex applications, such as decentralized financial instruments or autonomous organizations, are not feasible on the Bitcoin network. To overcome these limitations, the concept of programmable blockchains emerged, most notably with the introduction of Ethereum in 2015, see Buterin [6]. Ethereum extends the blockchain model by embedding a Turing-complete virtual machine, the Ethereum Virtual Machine (EVM), which enables the execution of code on-chain. This innovation allows developers to write smart contracts—autonomous programs that run exactly as programmed—opening the door to a wide range of decentralized applications (dApps) across finance, insurance, governance or digital ownership, some concrete projects are mentionned in Table 1.
@@ 3.1 Parametric insurance contract over time
A parametric insurance contract pays a predefined compensation $l$ if some observable quantity $(Q_{t})_{t\geq 0}$ reaches a threshold $\overline{Q}$ at some time $T$. The quantity $Q:=(Q_{t})_{t\geq 0}$ is a stochastic process defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. A compensation is paid if the $\mathcal{F}_{T}$-measurable event: $\{Q_{T}>\overline{Q}\}$ occurs. A parametric insurance contract is represented as a 5-tuple $(S,T,Q,\overline{Q},l)$, where $S$ denotes the time at which the contract is underwritten. We outline in Remark 3.1 and Remark 3.2 two extensions associated to the time of the triggering event and the amount compensated respectively.
@@
When the trigger is based on weather condition, the insurance coverage usually concerns a time interval. The triggering event definition must be tailored to the meteorological quantity we monitor. We consider daily rainfall as an example later in the article and so a typical event would be
@@
It is more convenient to treat $Q$ as a discrete time stochastic process or a time serie in that case. If we consider windspeed then time $T$ should be replaced by a time interval $(T_{-},T_{+})$ and the triggering event then becomes
$\{Q_{t}>\overline{Q}\text{, for some }t\in(T_{-},T_{+})\},\text{ or }\{Q_{t}>\overline{Q}\text{, for all }t\in(T_{-},T_{+})\},$
and $Q$ is a continuous time stochastic process.
@@
$\xi_{t}=\begin{cases}0,&\text{if }t<T,\\ 1,&\text{if }t\geq T\text{ and }Q_{t}\leq\overline{Q},\\ 2,&\text{if }t\geq T\text{ and }Q_{t}>\overline{Q}.\\ \end{cases}$ (1)
@@
[…]
The smart contract may be forced to file for bankruptcy if the surplus falls below the minimum capital requirement at any point between $S$ and $T$. If this occurs, the remaining active policies are cancelled, and the premiums are refunded to the policyholders. This possibility is described in Section 4.3.4.
The pure premium associated with a contract $(S,T,Q,\overline{Q},l)$ is given by $\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})\cdot l$, where $\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})$ is the probability of the event $\{Q_{T}>\overline{Q}\}$ estimated at time $S$. The filtration $(\mathcal{F}_{t})_{t\geq 0}$ gathers all the available information at time $t\leq T$ relevant to the occurrence of event $\{Q_{T}>\overline{Q}\}$ or equivalently the study of the quantity $(Q_{t})_{t\geq 0}$. For weather-linked events, the quantity $Q_{T}$ may be forecasted based on the available information at time $t<T$ using weather forecast models. Concretely, the filtration $(\mathcal{F}_{t})_{t\geq 0}$ contains both observations and predictions available at time $t\geq 0$ and this must be accounted for in our actuarial pricing framework. Constraints on $S$ can be imposed to mitigate the risk of adverse selection as outlined in Remark 3.3.
@@
From a practical standpoint, we can assume the existence of a time $S^{\ast}=T-\Delta<T$ before which the triggering event cannot be reliably predicted. Therefore, the estimation of the probability of $\{Q_{T}>\overline{Q}\}$ relies solely on historical, observed data. If policyholders have access to a more accurate predictive model than that of the insurers, it could lead to adverse selection. A straightforward approach to mitigate this risk is to stipulate that a policy $(S,T,Q,\overline{Q},l)$ cannot be underwritten if $S>S^{\ast}$. We can assume that:
$\theta=\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{s}),\text{ for any }s\leq S^{\ast}.$
We consider as a use case a parametric insurance coverage against rain episodes. An illustration is provided in 1 with a concrete estimation of the probability $\theta$.
@@
where $N_{T}$ is a Poisson random variable $\mathsf{Poisson}(\lambda_{T})$ and the $U_{k}$ are iid gamma random variables $\mathsf{Gamma}(\alpha_{T},\beta_{T})$ independent from $N_{T}$. This model was proposed by Dunn [11] with $N_{T}$ corresponding to the number of rain episodes during a day and the $U_{k}$’s characterize their intensity. We assume that the parameters of the model $\lambda_{T}$, $\alpha_{T}$ and $\beta_{T}$ simply depend on the month associated with the date $T$. We calibrate this model using daily data from the years $2000$ to $2023$ provided by Météo France and retrieved from the data.gouv.fr. The time series is shown in Figure 1.
@@
$\widehat{\alpha}=\frac{\mu^{2}}{\widehat{\lambda}\sigma^{2}-\mu^{2}}\text{ and}\widehat{\beta}=\frac{\mu}{\widehat{\lambda}\widehat{\alpha}}.$
This simple method is described for instance in Goffard et al. [15]. We have a compound Poisson-Gamma model for the STRASBOURG-ENTZHEIM station for each month; the estimations of the parameters allow us to estimate the number of rain episodes and their intensity for each month as shown in Figure 2.
@@
The climate in Strasbourg is "semi-continental". In terms of precipitation, it is characterized by low precipitation levels during the winter months (replaced by snow) and few but intense rain episodes during the summer months. The model then provides us with the probability $\theta$ associated with the threshold $\overline{Q}$ for each month, see Figure 3.
@@
The pricing of our parametric insurance contract relies on a piecewise constant function $T\mapsto\theta(T)$. From a practical perspective, dealing with dates in Solidity (the programming language of the Ethereum blockchain) is not straightforward; furthermore, having to handle 12 clusters of dates is tedious. Our solution consists of writing $T$ as an integer number between 1 and 365 and approximating the function $T\mapsto\theta(T)$ through a polynomial. In our smart contract, we only consider one threshold with $\overline{Q}=5$. The polynomial approximation of degree 4 is provided in Figure 4.
@@
Suppose that $S\leq S^{\ast}$, where $S^{\ast}$ has been defined in Remark 3.3. The commercial premium is then given by

### v3

@@ 1 Introduction
A blockchain is a decentralized ledger maintained by consensus within a peer-to-peer network. The nodes in this network agree on a shared history of data by executing an algorithm known as a consensus protocol. The Bitcoin blockchain, as described by Nakamoto [24], simply records transactions between bitcoin users. In contrast, the Ethereum blockchain, introduced by Buterin [6], is also equipped with a compiler to execute pieces of code called smart contract. Ethereum is designed to operate as a "world computer," enabling developers to deploy decentralized applications (dApps) that users can interact with directly.
Through these decentralized applications, Ethereum extends the capabilities of blockchains beyond simple fund transfers, enabling more complex financial operations and fostering the emergence of a new financial ecosystem known as decentralized finance (DeFi), as described by Schär [27]. Traditional finance relies on trusted third parties and centralized databases, which involve intermediary costs, lack transparency, operate only during limited hours, and are susceptible to cyberattacks. In contrast, distributed ledgers are publicly accessible, replicated in multiple locations, and function continuously, 24/7. Furthermore, the code that governs decentralized applications is auditable, improving transparency. For further discussion, we refer the reader to Lipton and Treccani [20, Chapter 2].
@@
In recent years, parametric insurance schemes have emerged in mass-market products, such as flight delay insurance offered by China Southern Airlines, which provides payouts of up to 300 yuan (approximately US\$40) for delays of three hours or more. Note that death benefits in traditional insurance often consist of a fixed amount that varies according to limited number of parameters, such as the number of children or marital status. This forms a prominent example of a parametric insurance contract that is widely accepted and commercialized.
The main downside of parametric insurance is basis risk, where the insured’s losses may not match the coverage amount, or the parameter may not trigger. Effective structuring and pricing require understanding the policyholder’s exact risks and choosing the right parameter. The availability of finer granularity data and advances in predictive modeling can mitigate basis risk. Hybrid insurance solutions that combine parametric and standard indemnity mechanisms can also be effective, see for instance Lopez and Nkameni [21]. An initial payment following the parametric trigger can provide immediate relief, and this amount can later be deducted from the compensation resulting from the traditional claims assessment process. The policyholder has the option to either maintain parametric insurance or purchase an additional traditional insurance contract to cover basis risk.
@@
We present a smart contract on the Ethereum blockchain to manage a portfolio of parametric insurance policies. We have developed an actuarial framework suited for weather-linked triggering events, specifically monitoring daily rainfall. This framework establishes ground rules for underwriting, rate-making, and managing risk. Our smart contract is designed to operate with minimal human interference, maximizing automation. Every interaction with the smart contract incurs a cost, as any transaction submitted to the blockchain requires a fee to be processed by the network of nodes. Therefore, optimizing the number of transactions is crucial for reducing operational costs. Actuarial methods for calculating risk capital are often too complex to be directly implemented in a high-level language like Solidity, which, for instance, only supports integer-based arithmetic. Additionally, deploying a smart contract on the blockchain entails costs that grow with its complexity, emphasizing the need for simplicity in design. One of our key contributions is the simplification of these computations through justified approximations. In particular, we propose a straightforward recursive formula to determine the solvency capital required to adequately cover the risks of a parametric insurance portfolio.
@@
The smart contract structure is inspired by the discussion in Cousaert et al. [10], from which we adopt the role distribution and the "tokenization" mechanism of risk. The smart contract owner writes the code and sets the initial values for parameters such as the premium loading and the risk level required for risk capital allocation. Investors participate by locking funds in the smart contract, in exchange for which they receive protocol tokens. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. The contribution of each investor is reflected in their token holdings. The value of the token, or equivalently its exchange rate against ETH, fluctuates over time depending on whether compensation are being paid to policyholders. Conversely, the token value increases if no compensation is paid at settlement. Token ownership enables governance of the insurance protocol. The contract owner may update the smart contract parameters based on recommendations from a board composed of token holders. This governance model transforms the smart contract into a Decentralized Autonomous Organization (DAO). The tokenization of risk on the blockchain can be viewed as a novel way to design insurance-linked securities, as outlined in Barrieu and Albertini [4]. By enabling policyholders to become investors, our smart contract creates a collaborative insurance scheme, akin to those described in Feng [12]. Investors, referred to as Surplus Providers (SPs), are analogous to liquidity providers in decentralized exchanges, which serve as blockchain-based trading venues for crypto assets; see, for instance, Mohan [23].
@@
This paper aims to show the feasibility of deploying a beta version of such a smart contract, where certain complexities are intentionally simplified. We also identify the main limitations of this simplified framework and provide the research and actuarial communities with an open-source beta version of the smart contract. This prototype serves as a foundation for developing generalizations and improved versions that address these limitations.
The rest of the paper is organized as follows. Section˜2 provides a brief overview on blockchain technology. Section˜3 presents an actuarial framework for managing parametric insurance contract portfolios. Section˜4 focuses on the design of the smart contract. We introduce the participants, including the contract owner, surplus providers, and policyholders, in Section˜4.1. A smart contract is characterized by state variables, which we describe in Section˜4.2. Examples of state variables include the smart contract’s balance and the tokens held by each participant. Users interact with the smart contract by invoking functions, and the possible actions and their impacts on the system’s state are detailed in Section˜4.3. Section˜5 demonstrates how the smart contract operates through an event-driven scenario, showcasing all its functions and culminating in the liquidation and reset of the smart contract. The main limitations of this beta version of the smart contract are discussed in Section˜6, along with potential ideas for addressing them.
@@ 2 Blockchain technology and smart contracts
However, the Bitcoin blockchain was designed with a narrow scope: it enables peer-to-peer transfer of value but offers very limited programmability. As a result, more complex applications, such as decentralized financial instruments or autonomous organizations, are not feasible on the Bitcoin network. To overcome these limitations, the concept of programmable blockchains emerged, most notably with the introduction of Ethereum in 2015, see Buterin [6]. Ethereum extends the blockchain model by embedding a Turing-complete virtual machine, the Ethereum Virtual Machine (EVM), which enables the execution of code on-chain. This innovation allows developers to write smart contracts—autonomous programs that run exactly as programmed—opening the door to a wide range of decentralized applications (dApps) across finance, insurance, governance or digital ownership, some concrete projects are mentionned in Table˜1.
@@ 3 Parametric insurance portfolio management framework
This section presents an actuarial framework for parametric insurance tailored to an implementation on the Ethereum blockchain. While this framework supports our analysis, we emphasize that it is not the primary contribution of our work. We further acknowledge that it could be extended or refined to accommodate a broader class of insurance solutions.
@@ 3.1 Parametric insurance contract over time
A parametric insurance contract pays a predefined compensation $l$ if some observable quantity $(Q_{t})_{t\geq 0}$ reaches a threshold $\overline{Q}$ at some time $T$. The quantity $Q:=(Q_{t})_{t\geq 0}$ is a stochastic process defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. A compensation is paid if the $\mathcal{F}_{T}$-measurable event: $\{Q_{T}>\overline{Q}\}$ occurs. A parametric insurance contract is represented as a 5-tuple $(S,T,Q,\overline{Q},l)$, where $S$ denotes the time at which the contract is underwritten. We outline in Remark˜3.1 and Remark˜3.2 two extensions associated to the time of the triggering event and the amount compensated respectively.
@@
When the trigger is based on weather condition, the insurance coverage usually concerns a time interval. The triggering event definition must be suited to the meteorological quantity we monitor. We consider daily rainfalls as an example later in the article and so a typical event would be
@@
We therefore consider $Q$ as a discrete time stochastic process or a time serie corresponding to the amount rain that fell on a specific day. A continuous time model would be better suited for other situations like flood risk for example. The trigger would be based on the cumulated rain over a rolling window of several days. If we consider windspeed then time $T$ should be replaced by a time interval $(T_{-},T_{+})$ and the triggering event then becomes
$\left\{\underset{t\in(T_{-},T_{+})}{\max}Q_{t}>\overline{Q}\right\},\text{ or }\left\{\underset{t\in(T_{-},T_{+})}{\min}Q_{t}>\overline{Q}\right\},$
where $Q$ is a continuous time stochastic process. Our approach could be extended to accomodate such refinements.
@@
[…]
@@
$0=\text{"open"},\quad 1=\text{"closed without compensation"},\text{ and}\quad 2=\text{"closed with compensation"}.$
The smart contract may be forced to file for bankruptcy if the surplus falls below the minimum capital requirement at any point between $S$ and $T$. If this occurs, the remaining active policies are cancelled, and the premiums are refunded to the policyholders. This possibility implies to add a fourth state
$3=\text{"cancelled"},$
which is covered in Section˜4.3.4.
The pure premium associated with a contract $(S,T,Q,\overline{Q},l)$ is given by
$\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})\cdot l.$
The probability of the event $\{Q_{T}>\overline{Q}\}$ estimated at some time $t\in[0,T)$ is a stochastic process defined by
$\theta_{t}=\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{t})\text{ for }t\in[0,T).$ (2)
The filtration $(\mathcal{F}_{t})_{t\geq 0}$ gathers all the available information at time $t\leq T$ relevant to the occurrence of event $\{Q_{T}>\overline{Q}\}$ or equivalently the study of the quantity $(Q_{t})_{t\geq 0}$. For weather-linked events, the quantity $Q_{T}$ may be forecasted based on the available information at time $t<T$ using weather forecast models. Concretely, the filtration $(\mathcal{F}_{t})_{t\geq 0}$ contains both observations and predictions available at time $t\geq 0$ and this must be accounted for in our actuarial pricing framework. Constraints on $S$ can be imposed to mitigate the risk of adverse selection as outlined in Remark˜3.3.
@@
From a practical standpoint, we can assume the existence of a time $S^{\ast}=T-\Delta<T$ before which the triggering event cannot be reliably anticipated. Therefore, the estimation of the probability of $\{Q_{T}>\overline{Q}\}$ relies solely on historical, observed data. If policyholders have access to a more accurate predictive model than that of the insurers, it could lead to adverse selection. A straightforward approach to mitigate this risk is to stipulate that a policy $(S,T,Q,\overline{Q},l)$ cannot be underwritten if $S>S^{\ast}$. We can assume that:
$\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{s})=\theta,\text{ for any }s\leq S^{\ast}.$
We consider as a use case a parametric insurance coverage against rain episodes. An illustration is provided in ˜1 with a concrete estimation of the probability $\theta$.
@@
where $N_{T}$ is a Poisson random variable $\mathsf{Poisson}(\lambda_{T})$ and the $U_{k}$ are iid gamma random variables $\mathsf{Gamma}(\alpha_{T},\beta_{T})$ independent from $N_{T}$. This model was proposed by Dunn [11] with $N_{T}$ corresponding to the number of rain episodes during a day and the $U_{k}$’s characterize their intensity. We assume that the parameters of the model $\lambda_{T}$, $\alpha_{T}$ and $\beta_{T}$ simply depend on the month associated with the date $T$. We calibrate this model using daily data from the years $2000$ to $2023$ provided by Météo France and retrieved from the data.gouv.fr. The time series is shown in Figure˜1.
@@
$\widehat{\alpha}=\frac{\mu^{2}}{\widehat{\lambda}\sigma^{2}-\mu^{2}}\text{ and }\widehat{\beta}=\frac{\mu}{\widehat{\lambda}\widehat{\alpha}}$,
where $\mu$ and $\sigma$ denote the empirical mean and variance of the daily precipitations for rainy days. This simple method is described for instance in Goffard et al. [15]. We have a compound Poisson-Gamma model for the STRASBOURG-ENTZHEIM station for each month; the estimations of the parameters allow us to estimate the number of rain episodes and their intensity for each month as shown in Figure˜2.
@@
The climate in Strasbourg is "semi-continental". In terms of precipitation, it is characterized by low precipitation levels during the winter months (replaced by snow) and few but intense rain episodes during the summer months. The model then provides us with the probability $\theta$ associated with the threshold $\overline{Q}$ for each month, see Figure˜3.
@@
The pricing of our parametric insurance contract relies on a piecewise constant function $T\mapsto\theta(T)$. From a practical perspective, dealing with dates in Solidity (the programming language of the Ethereum blockchain) is not straightforward; furthermore, having to handle 12 clusters of dates is tedious. Our solution consists of writing $T$ as an integer number between 1 and 365 and approximating the function $T\mapsto\theta(T)$ through a polynomial. In our smart contract, we only consider one threshold with $\overline{Q}=5$. The polynomial approximation of degree 4 is provided in Figure˜4.
@@
Suppose that $S\leq S^{\ast}$, where $S^{\ast}$ has been defined in Remark˜3.3. The commercial premium is then given by
@@
Remark 3.5.
An insurance coverage $(S,T,Q,\overline{Q},l)$ cannot be purchased after time $S^{\ast}$. To further mitigate the risk of adverse selection, the safety loading could also be adjusted to incentivize early purchases—specifically, by offering a premium discount to customers who buy the coverage well before $S^{\ast}$.